# FAB-PPI: Frequentist, Assisted by Bayes, Prediction-Powered Inference

### v1

@@ 4.1 Bayes-assisted PPI Estimators
for the rectifier $\Delta_{\theta}$. By Tweedie’s formula , the above estimator is the posterior mean of the mean parameter of a Gaussian likelihood model under the prior $\pi_{0}$. Note however that we do not assume here that $\widehat{\Delta}_{\theta}$ is normally distributed for a fixed $n$.
@@ 4.2 FAB-PPI Confidence Regions
  Set $\widehat{\lambda}=1$ (FAB-PPI) or estimate $\widehat{\lambda}$ from data (FAB-PPI++) as in Angelopoulos et al. (2023a).
@@
     $(\widehat{\sigma}^{f}_{\theta})^{2}\leftarrow\frac{1}{N-1}\sum_{i=1}^{N}\left(\mathcal{L}_{\theta}^{\prime}(\widetilde{X}_{i},f(\widetilde{X}_{i}))-\widehat{m}_{\theta}\right)^{2}$
@@
     $\mathcal{T}_{\alpha-\delta}(\widehat{m}_{\theta})\leftarrow\left(\widehat{m}_{\theta}\pm\widehat{\sigma}^{f}_{\theta}z_{1-(\alpha-\delta)/2}\right)$
@@ 4.3 Choosing the prior
FAB-PPI is motivated by applications in which the PPI predictor $f$ is expected to be generally accurate, as measured by the rectifier $\Delta_{\theta}$. Such a property may be encoded in $\pi_{0}(\Delta_{\theta};\tau_{n})$ by choosing a prior that concentrates around zero. As mentioned in Section 3.3, the FAB construction of $\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta})$ will exhibit smaller volume compared to the classical CI, and hence result in downstream efficiency gains over standard PPI, if the true rectifier $\Delta_{\theta}$ is likely under $\pi_{0}$. In particular, the prior scale $\tau_{n}$ controls the size of the potential efficiency gains and losses of FAB-PPI over PPI: the smaller $\tau_{n}$, the more the resulting CR will shrink (resp. grow) when $\Delta_{\theta}\approx 0$ (resp. $|\Delta_{\theta}|\gg 0$). Experimentally, we find that the choice $\tau_{n}=\widehat{\sigma}_{\theta}$ results in a parameter-free approach that strikes a good compromise. More general choices of $\tau_{n}$ are briefly mentioned in Section 6.
@@
[…]
@@ 4.4 Theoretical properties
For $\alpha\in(0,1)$, let $\mathcal{C}_{\alpha}^{\mathrm{\texttt{FABPP}}}$ be the FAB-PPI confidence region (21) under the Gaussian prior (22) or the horseshoe prior (23). Then, under 3.1,
$\liminf_{\min(n,N)\to\infty}\Pr(\theta^{*}\in\mathcal{C}_{\alpha}^{\mathrm{\texttt{FABPP}}})\geq 1-\alpha.$
The proof of Theorem 4.1 crucially relies on showing exact asymptotic coverage of the FAB CR $\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta})$. While the latter holds for both priors introduced in the previous sections, the two limits are very different. In particular, as discussed in Remark S3.5, the volume of $\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta})$ vanishes asymptotically under $\pi_{\text{HS}}$, while it does not under $\pi_{\text{N}}$.
@@
For $\alpha\in(0,1)$, let $\mathcal{C}_{\alpha}^{\mathrm{\texttt{FABPP}}}$ and $\mathcal{C}_{\alpha}^{\mathrm{\texttt{PP}}}$ denote, respectively, the FAB-PPI confidence region (21) under the horseshoe prior (23) and the standard PPI confidence region (8), both viewed as functions of $\widehat{\Delta}_{\theta}$. If $|\widehat{\Delta}_{\theta}|\gg 0$, then
@@
Overall, Remark S3.5 and Proposition 4.2 provide strong support for preferring $\pi_{\text{HS}}$ over $\pi_{\text{N}}$ within the FAB-PPI framework.
@@ 4.5 FAB-PPI for mean estimation
where $\widehat{\sigma}^{2}$ is an estimator of $\mathrm{var}(\widehat{\Delta})$, $\widehat{\theta}$ is the PPI estimator corresponding to $\widehat{\Delta}$, and $\widehat{\lambda}$ is set either to one (PPI) or (14) (PPI++). In both cases, the estimator $\widehat{\theta}^{\mathrm{\texttt{FABPP}}}$ takes the form
Classic Estimator + PPI correction + Bayes $\text{\normalsize Classic Estimator + PPI correction + {Bayes correction}},$
@@ 5 Experiments
We compare FAB-PPI and power-tuned FAB-PPI (FAB-PPI++) to classical inference, PPI and power-tuned PPI (PPI++) on both synthetic and real estimation problems. For FAB-PPI, we use (HS) and (N) to indicate the use of the horseshoe and Gaussian priors defined in Section 4.3. As already mentioned, PPI is motivated by settings in which labelled data are scarce, while unlabelled data are abundant. Moreover, the application of FAB to PPI specifically targets the estimation of the rectifier $\Delta_{\theta}$. For these reasons, we choose to focus on cases where $N\gg n$ is large enough to rule out any uncertainty on the measure of fit $m_{\theta}$, which we estimate using the sample mean $\widehat{m}_{\theta}$ (9). As a result of this, given a $1-\delta$ confidence interval (FAB or not) $\mathcal{R}_{\delta}$ for $\Delta_{\theta}$, the corresponding $1-\alpha$ CI for $\theta^{\star}$ is obtained simply by setting $\delta=\alpha$ and shifting $\mathcal{R}_{\delta}$ by $\widehat{m}_{\theta}$. This simplification allows us to evaluate the direct effect of FAB on the procedure, eliminating concerns about the loss of tightness in the CI on $\theta^{*}$ due to the Minkowski sum in Equation 21. In all experiments, we check empirically that $N$ is large enough to make this assumption by monitoring the coverage of the resulting intervals against both the nominal level $1-\alpha$ and the coverage of PPI intervals that also consider the uncertainty on $m_{\theta}$ (denoted with PPI (full) and PPI++ (full) in Appendix S6).
@@ 5.1 Synthetic data
The simulated experiments below have a common structure. We sample two datasets, $n$ labelled observations $\{(X_{i},Y_{i})\}_{i=1}^{n}$ iid from $\mathbb{P}$ and $N$ unlabelled observations $\{\widetilde{X}_{i}\}_{i=1}^{N}$ iid from $\mathbb{P}_{X}$. We use a prediction rule $f$ to obtain predictions $\{f(X_{i})\}_{i=1}^{n}$ and $\{f(\widetilde{X}_{i})\}_{i=1}^{N}$. We apply the different procedures to obtain estimates and $1-\alpha$ confidence regions for the mean $\theta^{*}=\mathbb{E}[Y]$. For all experiments, we set $\alpha=0.1$ and report the average mean squared error (MSE), interval volume, and coverage over $1000$ repetitions.
@@
We sample $X_{i}\overset{iid}{\sim}\mathcal{N}(0,1)$ and $Y_{i}=X_{i}+\epsilon_{i}$ with $\epsilon_{i}\overset{iid}{\sim}\mathcal{N}(0,1)$, so that $\theta^{*}=\mathbb{E}[Y]=0$. The prediction rule is defined as $f(X_{i})=X_{i}+\gamma$, where $\gamma\in\mathbb{R}$. For this choice, the bias of $f$ is controlled by $\gamma$, since $\text{MSE}(f)=\gamma^{2}+1$. For this experiment, we assume that $N$ is infinite, set $n=200$, and vary $\gamma$ between $-1.5$ and $1.5$. Figure 1 shows the average interval volume as a function of $\gamma$ for classical inference, PPI++, and FAB-PPI++ with both a horseshoe and a Gaussian prior.
@@
Results for the non power-tuned methods, as well as MSE and coverage plots, are reported in Figure S6. Except for the version with the Gaussian prior, all the PPI procedures outperform classical inference for every bias level $\gamma$, but the behaviour exhibited by PPI++ deserves attention, as its CI volume is approximately constant across values of $\gamma$. This is due to the fact that, since $N$ is taken to be infinite and $n$ is fairly large, $\widehat{\lambda}\approx\mathrm{cov}(Y,f(X))=1$ and the rectifier is accurately estimated with similar variance across all values of $\gamma$. On the other hand, the CI volume for the FAB-PPI methods varies greatly with $\gamma$. When the bias is small ($\gamma\simeq 0$), the observed rectifier has a value close to 00, leading to smaller CIs. As the bias increases, the volume of the confidence intervals grows, until it surpasses that of the PPI intervals. At this point, the two FAB-PPI procedures behave differently: the volume of the Gaussian intervals grows unbounded, whereas the horseshoe intervals eventually revert to the PPI ones. This example clearly shows that FAB-PPI with a horseshoe prior allows to obtain smaller CIs when the predictions are good, while ensuring robustness as the quality of the predictions decreases (Proposition 4.2).
@@
We consider the mean estimation example of Angelopoulos et al. (2023a, §7.1.1), which does not involve any covariate $X$. We sample $Y_{i}\overset{iid}{\sim}\mathcal{N}(0,1)$, so that $\theta^{*}=\mathbb{E}[Y]=0$. The prediction rule is defined as $f(X_{i})=Y_{i}+\sigma_{Y}\epsilon_{i}$, where $\epsilon_{i}\overset{iid}{\sim}\mathcal{N}(0,1)$ and $\sigma_{Y}$ is successively set to $0.1$, $1$, and $2$. For this experiment, we set $N=10^{6}$ and vary $n$ from $100$ to $1000$. Figure 2 shows the average interval volume as a function of $n$ for the different methods as the noise level $\sigma_{Y}$ varies, while similar plots for the MSE and coverage are reported in Figure S7.
@@
In this case, the effect of power tuning matches the observations of Angelopoulos et al. (2023a): as the noise level increases, $\hat{\lambda}$ decreases and less weight is given to the predicted labels. When the noise is small, all PPI procedures perform similarly, and much better than classical inference. When the noise is large, the power-tuned procedures perform similar to or better than classical inference, whereas the non-tuned alternatives lose ground. At the intermediate noise level, the power-tuned methods clearly outperform the other baselines. Crucially, FAB-PPI outperforms the PPI counterpart at all noise levels, with FAB-PPI++ being the best performer overall. This is because, in this setting, while predictions exhibit increasing variance with $\sigma_{Y}$, they remain unbiased. As a result of this, regardless of the value of $\lambda$ used, any additional shrinkage performed on the rectifier by FAB-PPI is beneficial. This example shows that FAB-PPI++ retains the benefits of power tuning, while also taking advantage of the adaptive shrinkage provided by the FAB procedure.
@@ 5.2 Real data
We consider several estimation experiments using the datasets presented in Angelopoulos et al. (2023b) and briefly described in Section S5.1. Each dataset comes with covariate/label/prediction triples $\{X_{i},Y_{i},f(X_{i})\}_{i=1}^{N}$, which we randomly split into two subsets with $n$ labelled and $N-n$ unlabelled observations, for varying values of $n$. For all experiments and methods, we report the average estimation MSE, CI volume and coverage across multiple repetitions.
We begin with three mean estimation experiments, involving the alphafold, forest, and galaxies datasets, and one logistic regression experiment on the healthcare dataset. Figure 3 shows the results for classical inference, PPI++, and FAB-PPI++ applied to the mean estimation datasets. The mean estimation results for the non power-tuned methods are reported in Figure S8, whereas the ones for the logistic regression experiment are reported in Figure S11. In all cases, FAB-PPI/FAB-PPI++ outperform classical inference and the corresponding PPI methods, both in terms of MSE and CI volume, while achieving comparable coverage. These examples suggest that the quality of the predictions of existing machine learning models on several real datasets may fall into the regime where the adaptive shrinkage provided by the FAB framework leads to a further improvement over standard PPI. In these settings, as the predictions are good, FAB-PPI under the horseshoe and Gaussian priors exhibit similar gains, as already seen in Figure 1.
However, the same is not true in the presence of bad predictions. For instance, Figure S12 shows the results of a quantile estimation experiment on the genes dataset, where predictions are heavily biased. In this case, the behaviour of the FAB-PPI methods under the horseshoe and Gaussian priors differs significantly: the former matches the performance of the PPI methods, which outperform classical inference, whereas the latter leads to much larger MSE and CIs. As previously discussed, such desirable behaviour of FAB-PPI under the horseshoe prior is due to its robustness against large bias levels (Proposition 4.2). Similarly, Figure S13 reports the results of a linear regression experiment on the census dataset. For one of the two parameters considered (panel (a)), FAB-PPI underperforms the alternatives under both priors for small $n$. However, as $n$ increases, the performance under the horseshoe prior improves and eventually matches that of the PPI methods, while the Gaussian prior does not. This example shows another facet of the horseshoe’s robustness: even for moderate bias levels, as the available labelled sample size grows, disagreements between the prior and the data become apparent (i.e. $\mathrm{var}(\widehat{\Delta}_{\theta})$ decreases), eventually leading Proposition 4.2 to kick in.

### v2

@@ 4.1 Bayes-Assisted PPI Estimators
for the rectifier $\Delta_{\theta}$. By Tweedie’s formula (Efron, 2011), the above estimator is the posterior mean of the mean parameter of a Gaussian likelihood model under the prior $\pi_{0}$. Note however that we do not assume here that $\widehat{\Delta}_{\theta}$ is normally distributed for a fixed $n$.
@@ 4.2 FAB-PPI Confidence Regions
 Set $\widehat{\lambda}=1$ (FAB-PPI) or estimate $\widehat{\lambda}$ from data (FAB-PPI++) as in Angelopoulos et al. (2023b).
@@
  $\widehat{\sigma}_{m}^{2}\leftarrow\frac{1}{N-1}\sum_{i=1}^{N}\left(\mathcal{L}_{\theta}^{\prime}(\widetilde{X}_{i},f(\widetilde{X}_{i}))-\widehat{m}_{\theta}\right)^{2}$
@@
  $\mathcal{T}_{\alpha-\delta}(\widehat{m}_{\theta})\leftarrow\left(\widehat{m}_{\theta}\pm\frac{\widehat{\sigma}_{m}}{\sqrt{N}}z_{1-(\alpha-\delta)/2}\right)$
@@ 4.3 Choosing the Prior
FAB-PPI is motivated by applications in which the PPI predictor $f$ is expected to be generally accurate, as measured by the rectifier $\Delta_{\theta}$. Such a property may be encoded in $\pi_{0}(\Delta_{\theta};\tau_{n})$ by choosing a prior that concentrates around zero. As mentioned in Section 3.3, the FAB construction of $\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta})$ will exhibit smaller volume compared to the classical CI, and hence result in downstream efficiency gains over standard PPI, if the true rectifier $\Delta_{\theta}$ is likely under $\pi_{0}$. In particular, the prior scale $\tau_{n}$ controls the size of the potential efficiency gains and losses of FAB-PPI over PPI: the smaller $\tau_{n}$, the more the resulting CR will shrink (resp. grow) when $\Delta_{\theta}\simeq 0$ (resp. $|\Delta_{\theta}|\gg 0$). Experimentally, we find that the choice $\tau_{n}=\widehat{\sigma}_{\theta}$ results in a parameter-free approach that strikes a good compromise. More general choices of $\tau_{n}$ are briefly mentioned in Section 6.
@@
$\pi_{\text{HS}}(\Delta_{\theta};\widehat{\sigma}_{\theta})=\int_{0}^{\infty}\mathcal{N}(\Delta_{\theta};0,\nu^{2}\widehat{\sigma}^{2}_{\theta})C^{+}(\nu;0,1)d\nu,$ (23)
@@ 4.4 Theoretical Properties
For $\alpha\in(0,1)$, let $\mathcal{C}_{\alpha}^{\mathrm{\texttt{FABPP}}}$ be the FAB-PPI confidence region (21) under the Gaussian prior (22) or the horseshoe prior (23). Then, under Section 3.2,
$\liminf_{\min(n,N)\to\infty}\Pr(\theta^{\star}\in\mathcal{C}_{\alpha}^{\mathrm{\texttt{FABPP}}})\geq 1-\alpha.$
The proof of Theorem 4.1 crucially relies on showing exact asymptotic coverage of the FAB CR $\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta})$. While the latter holds for both priors introduced in the previous sections, the two limits behave very differently. In particular, as discussed in Section S3.1, the volume of $\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta})$ vanishes asymptotically under $\pi_{\text{HS}}$, while it does not under $\pi_{\text{N}}$.
@@
For $\alpha\in(0,1)$, let $\mathcal{C}_{\alpha}^{\mathrm{\texttt{FABPP}}}$ and $\mathcal{C}_{\alpha}^{\mathrm{\texttt{PP}}}$ denote, respectively, the FAB-PPI confidence region (21) under the horseshoe prior (23) and the standard CLT-based PPI CI for $\theta$, both viewed as functions of $\widehat{\Delta}_{\theta}$. If $|\widehat{\Delta}_{\theta}|\gg 0$, then
@@
Overall, Section S3.1 and Section 4.4 provide strong support for preferring $\pi_{\text{HS}}$ over $\pi_{\text{N}}$ within the FAB-PPI framework.
@@ 4.5 FAB-PPI for Mean Estimation
[…]
$\text{\normalsize Classic Estimator + PPI correction + {Bayes correction}},$
@@ 5 Experiments
We compare FAB-PPI and power-tuned FAB-PPI (FAB-PPI++) to classical inference, PPI and power-tuned PPI (PPI++) on both synthetic and real estimation problems. For FAB-PPI, we use (HS) and (N) to indicate the use of the horseshoe and Gaussian priors defined in Section 4.3. As already mentioned, PPI is motivated by settings in which labelled data are scarce, while unlabelled data are abundant. Moreover, the application of FAB to PPI specifically targets the estimation of the rectifier $\Delta_{\theta}$. For these reasons, we choose to focus on cases where $N\gg n$ is large enough to rule out any uncertainty in the measure of fit $m_{\theta}$, which we estimate using the sample mean $\widehat{m}_{\theta}$ (9). As a result of this, given a $1-\delta$ confidence interval (FAB or not) $\mathcal{R}_{\delta}$ for $\Delta_{\theta}$, the corresponding $1-\alpha$ CI for $\theta^{\star}$ is obtained simply by setting $\delta=\alpha$ and shifting $\mathcal{R}_{\delta}$ by $\widehat{m}_{\theta}$. This simplification allows us to evaluate the direct effect of FAB on the procedure, eliminating concerns about the loss of tightness in the CI on $\theta^{\star}$ due to the Minkowski sum in Equation 21. In all experiments, we check empirically that $N$ is large enough to make this assumption by monitoring the coverage of the resulting intervals against both the nominal level $1-\alpha$ and the coverage of PPI intervals that also consider the uncertainty in $m_{\theta}$ (denoted with PPI (full) and PPI++ (full) in Appendix S6).
@@ 5.1 Synthetic Data
The simulated experiments below have a common structure. We sample two datasets, $n$ labelled observations $\{(X_{i},Y_{i})\}_{i=1}^{n}$ iid from $\mathbb{P}$ and $N$ unlabelled observations $\{\widetilde{X}_{i}\}_{i=1}^{N}$ iid from $\mathbb{P}_{X}$. We use a prediction rule $f$ to obtain predictions $\{f(X_{i})\}_{i=1}^{n}$ and $\{f(\widetilde{X}_{i})\}_{i=1}^{N}$. We apply the different procedures to obtain estimates and $1-\alpha$ confidence regions for the mean $\theta^{\star}=\mathbb{E}[Y]$. For all experiments, we set $\alpha=0.1$ and report the average mean squared error (MSE), interval volume, and coverage over $1000$ repetitions.
@@
We sample $X_{i}\overset{iid}{\sim}\mathcal{N}(0,1)$ and $Y_{i}=X_{i}+\epsilon_{i}$ with $\epsilon_{i}\overset{iid}{\sim}\mathcal{N}(0,1)$, so that $\theta^{\star}=\mathbb{E}[Y]=0$. The prediction rule is defined as $f(X_{i})=X_{i}+\gamma$, where $\gamma\in\mathbb{R}$. For this choice, the bias of $f$ is controlled by $\gamma$, since $\text{MSE}(f)=\gamma^{2}+1$. For this experiment, we assume that $N$ is infinite, set $n=200$, and vary $\gamma$ between 1.5-1.5 and $1.5$. Figure 1 shows the average interval volume as a function of $\gamma$ for classical inference, PPI++, and FAB-PPI++ with both a horseshoe and a Gaussian prior.
@@
Results for the non power-tuned methods, as well as MSE and coverage plots, are reported in Figure S6. Except for the version with the Gaussian prior, all the PPI procedures outperform classical inference for every bias level $\gamma$, but the behaviour exhibited by PPI++ deserves attention, as its CI volume is approximately constant across values of $\gamma$. This is due to the fact that, since $N$ is taken to be infinite and $n$ is fairly large, $\widehat{\lambda}\simeq\mathrm{cov}(Y,f(X))=1$ and the rectifier is accurately estimated with similar variance across all values of $\gamma$. On the other hand, the CI volume for the FAB-PPI methods varies greatly with $\gamma$. When the bias is small ($\gamma\simeq 0$), the observed rectifier has a value close to 0, leading to smaller CIs. As the bias increases, the volume of the confidence intervals grows, until it surpasses that of the PPI intervals. At this point, the two FAB-PPI procedures behave differently: the volume of the Gaussian intervals grows without bound, whereas the horseshoe intervals eventually revert to the PPI ones. This example clearly shows that FAB-PPI with a horseshoe prior allows to obtain smaller CIs when the predictions are good, while ensuring robustness as the quality of the predictions decreases (Section 4.4).
@@
We consider the mean estimation example of Angelopoulos et al. (2023b, §7.1.1), which does not involve any covariate $X$. We sample $Y_{i}\overset{iid}{\sim}\mathcal{N}(0,1)$, so that $\theta^{\star}=\mathbb{E}[Y]=0$. The prediction rule is defined as $f(X_{i})=Y_{i}+\sigma_{Y}\epsilon_{i}$, where $\epsilon_{i}\overset{iid}{\sim}\mathcal{N}(0,1)$ and $\sigma_{Y}$ is successively set to $0.1$, 11, and $2$. For this experiment, we set $N=10^{6}$ and vary $n$ from $100$ to $1000$. Figure 2 shows the average interval volume as a function of $n$ for the different methods as the noise level $\sigma_{Y}$ varies, while similar plots for the MSE and coverage are reported in Figure S7.
@@
In this case, the effect of power tuning matches the observations of Angelopoulos et al. (2023b): as the noise level increases, $\hat{\lambda}$ decreases and less weight is given to the predicted labels. When the noise is small, all PPI procedures perform similarly, and much better than classical inference. When the noise is large, the power-tuned procedures perform similarly to or better than classical inference, whereas the non-tuned alternatives lose ground. At the intermediate noise level, the power-tuned methods clearly outperform the other baselines. Crucially, FAB-PPI outperforms the PPI counterpart at all noise levels, with FAB-PPI++ being the best performer overall. This is because, in this setting, while predictions exhibit increasing variance with $\sigma_{Y}$, they remain unbiased. As a result of this, regardless of the value of $\lambda$ used, any additional shrinkage performed on the rectifier by FAB-PPI is beneficial. This example shows that FAB-PPI++ retains the benefits of power tuning, while also taking advantage of the adaptive shrinkage provided by the FAB procedure.
@@ 5.2 Real Data
We consider several estimation experiments using the datasets presented in Angelopoulos et al. (2023a) and briefly described in Section S5.1. Each dataset comes with covariate/label/prediction triples $\{X_{i},Y_{i},f(X_{i})\}_{i=1}^{N}$, which we randomly split into two subsets with $n$ labelled and NnN-n unlabelled observations, for varying values of $n$. For all experiments and methods, we report the average estimation MSE, CI volume and coverage across multiple repetitions.
We begin with four experiments, where the machine learning predictions provided are of high quality, and whose goals are as follows. Two of them are mean estimation tasks performed on the galaxies and forest datasets. The third one, performed on the alphafold dataset, is an odds ratio estimation task, for which the construction of confidence intervals also indirectly involves mean estimation as detailed in Section S5.1. The fourth one, involving the healthcare dataset, is a logistic regression task. Figure 3 shows the results for classical inference, PPI++, and FAB-PPI++ applied to the datasets involving mean estimation.
The mean estimation results for the non power-tuned methods are reported in Figure S8, whereas the ones for the logistic regression experiment are reported in Figure S11. In all cases, FAB-PPI/FAB-PPI++ outperform classical inference and the corresponding PPI methods, both in terms of MSE and CI volume, while achieving comparable coverage. These examples suggest that the quality of the predictions of existing machine learning models on several real datasets may fall into the regime where the adaptive shrinkage provided by the FAB framework leads to a further improvement over standard PPI. In these settings, as the predictions are good, FAB-PPI under the horseshoe and Gaussian priors exhibit similar gains, as already seen in Figure 1.
However, the same is not true in the presence of bad predictions. For instance, Figure S12 shows the results of a quantile estimation experiment on the genes dataset, where predictions are heavily biased. In this case, the behaviour of the FAB-PPI methods under the horseshoe and Gaussian priors differs significantly: the former matches the performance of the PPI methods, which outperform classical inference, whereas the latter leads to much larger MSE and CIs. As previously discussed, such desirable behaviour of FAB-PPI under the horseshoe prior is due to its robustness against large bias levels (Section 4.4). Similarly, Figure S13 reports the results of a linear regression experiment on the census dataset. For one of the two parameters considered (panel (a)), FAB-PPI underperforms the alternatives under both priors for small $n$. However, as $n$ increases, the performance under the horseshoe prior improves and eventually matches that of the PPI methods, while the Gaussian prior does not. This example shows another facet of the horseshoe’s robustness: even for moderate bias levels, as the available labelled sample size grows, disagreements between the prior and the data become apparent (i.e. $\mathrm{var}(\widehat{\Delta}_{\theta})$ decreases), eventually leading Section 4.4 to take effect.